the product of two numbers is 2. The sum of one number the reciprocal of the other is 6. Find the number?
Answer by checkley71(8403)

XY=2 OR X=2/Y
X+1/Y=6
2/Y+1/Y=6
3/Y=6
3=6Y
6Y=3
Y=3/6
Y=1/5 ANSWER.
X*1/2=2
X/2=2
X=282
X=4 ANSWER.
PROOF
4+1/(1/2)=6
4+1*2=6
6=6